# CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection

### v1

@@ 4.1 Experimental Setup
For safety benchmarks, both the refusal benchmark HarmBench and the overrefusal benchmark XSTest are included to evaluate not only the steering ability for rejection but also the model’s capability to identify the context of requests.
@@
Evaluation Metrics: For multiple-choice tasks (MMLU, MMLU-Pro, BBQ), exact match accuracy is used. For safety benchmarks, 1 - ASR (Attack Success Rate) is computed using a small refusal-detection language model. SimpleQA performance is measured using a small STS language model to match the expected answer, with more details in Appendix A.2.
@@
Pooling Strategy: Two pooling strategies are available for coefficient and correlation calculations: max-pooling and mean-pooling. For multi-token generation tasks, max-pooling consistently outperforms mean-pooling, as empirically demonstrated in Table 5, likely due to its better capture of peak feature activations relevant to task success. However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values. Applying these large coefficients to every generated token degrades performance, leading to the adoption of mean-pooling for reasoning tasks.
4.2 SER Comparison
This study aims to demonstrate that CorrSteer can improve benchmark performance while maintaining low side effects. Using the defined SER, CorrSteer’s SER is compared against fine-tuning for a question-answering dataset. Additionally, the SER is compared when allowing or enforcing negatively correlated features, supporting the claim of SAE’s positive-tinted behavior. Finally, the SER is compared when pooling on every token rather than the inference-time generated token.
4.3 Feature Interpretability Analysis
The analysis is enhanced by incorporating feature descriptions from Neuronpedia, providing semantic interpretations of selected features. Safe/unsafe tendency analysis is conducted, along with task-wise and feature-level interpretations. This interpretability analysis validates that our correlation-based selection identifies semantically meaningful and task-relevant features, supporting the causal hypothesis underlying our approach.
Table 1: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, GSM8K, SimpleQA, and XSTest on Gemma 2 2B. Results show accuracy (%) for all tasks.
Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
BBQ Ambig 59.10 62.38 66.65 62.08 -
BBQ Disambig 75.42 75.70 77.04 76.53 -
HarmBench 44.64 46.07 51.79 67.50 -
MMLU 52.23 52.82 55.83 56.32 55.85
MMLU-Pro 30.30 30.44 30.82 31.01 33.16
GSM8K 54.74 54.44 53.53 54.44 47.38
SimpleQA 3.63 3.76 3.96 3.80 -
XSTest 86.35 86.67 86.03 87.30 -
Table 2: Performance comparison between baseline and CorrSteer variants across BBQ, MMLU, MMLU-Pro, HarmBench, SimpleQA, and XSTest on LLaMA 3.1 8B. Results show accuracy (%) for all tasks.
Task Baseline CorrSteer-1 CorrSteer-P CorrSteer-A
@@ 4.3 Feature Interpretability Analysis
5 Results
5.1 Generation Benchmark Performance
Table 1 and Table 2 present comprehensive results across our evaluation benchmarks. CorrSteer demonstrates consistent improvements across most tasks, including question answering, bias mitigation, and safety benchmarks, with the exception of the GSM8K reasoning task. This pattern suggests that our method enhances model adherence to static task requirements without introducing dynamic behaviors.
Method Comparison
In most cases, CorrSteer-A and CorrSteer-P demonstrate the highest performance, with CorrSteer-P showing particular dominance in LLaMA 3.1 8B. This observation is attributed to the less disentangled nature of Llama Scope features from superposition, which necessitates more aggressive pruning.
Safety Benchmarks:
For HarmBench, selected features enhance the model’s ability to refuse harmful requests, achieving a 22.9% improvement. However, for XSTest, improvement is limited to 1% due to the overrefusal characteristics of the benchmark. This is an expected result due to the static nature of CorrSteer, which hinders the ability to clearly distinguish between benign and harmful requests.
Table 3: Safety Evaluation Rate (SER) analysis for CorrSteer variants across different benchmarks on Gemma 2 2B. SER values closer to 0 indicate better safety performance.
CorrSteer-1 CorrSteer-P CorrSteer-A Fine-tuning
Task SER neg pos SER neg pos SER neg pos SER neg pos
BBQ Ambig 0.000 0 658 0.000 0 1532 0.076 53 649 - - -
BBQ Disambig 0.167 45 59 0.153 111 164 0.257 65 112 - - -
HarmBench 0.250 2 6 0.143 4 24 0.043 3 67 - - -
MMLU 0.355 11 20 0.172 249 286 0.202 264 299 0.407 1108 1616
MMLU-Pro 0.421 8 11 0.423 30 41 0.419 39 54 0.461 357 418
GSM8K 0.556 20 16 0.674 31 15 0.516 63 59 0.647 213 116
SimpleQA 0.167 1 5 0.188 3 13 0.353 6 11 - - -
XSTest 0.333 7 10 0.520 7 10 0.467 14 5 - - -
Table 4: Safety Evaluation Rate (SER) analysis for CorrSteer variants on Llama models across different benchmarks. SER values closer to 0 indicate better safety performance.
CorrSteer-1 CorrSteer-P CorrSteer-A
@@ 5.1 Generation Benchmark Performance
Observations:
Results for both Gemma-2 2B and LLaMA 3.1 8B, presented in Table 1 and Table 2, demonstrate consistent patterns, with CorrSteer variants showing systematic improvements.
Multi-layer Superiority: Simultaneous multi-layer steering approaches, such as CorrSteer-A and CorrSteer-P, consistently outperform the single-layer CorrSteer-1 approach across all benchmarks, consistent with findings from other researchers (Liu et al., 2024; Zhao et al., 2025). This observation suggests that enabling features from different layers to compete globally yields superior task-relevant selections compared to layer-wise optimization.
Limited Factuality Impact: SimpleQA shows minimal improvement, confirming that CorrSteer enhances task adherence without introducing external factual knowledge. This is a desirable property, as it indicates the method improves model behavior rather than injecting information not present in the original model.
Comparison with Fine-tuning: CorrSteer demonstrates competitive performance while maintaining significantly lower side effect rates compared to fine-tuning. On MMLU, CorrSteer-A achieves higher accuracy (56.32%) than fine-tuning (55.85%) with substantially lower SER (0.202 vs 0.407). Similarly, on GSM8K, CorrSteer variants outperform fine-tuning in accuracy while maintaining lower side effect rates across all tasks.
5.2 Feature Analysis
Selected features demonstrate varying degrees of alignment with task requirements: while some benchmarks show consistent alignment through structured output features for multiple-choice tasks, refusal-related features for safety benchmarks, and task-specific semantic features for domain-specific evaluations, others exhibit less consistent patterns. The interpretability of selected features, validated through Neuronpedia descriptions, provides confidence in the semantic relevance of our correlation-based selection process. Notably, feature activation frequencies vary significantly across tasks, with performance improvements correlating with distinct activation patterns (Appendix 8). Analysis of selected features reveals semantically meaningful patterns aligned with task requirements:
Safety-related Features: HarmBench and BBQ demonstrate substantial improvements through selected safety-related features that enhance neutrality and refusal behavior. This interpretability analysis is covered in detail in Section 6.1.
5.3 SER Analysis
Table 3 presents the Safety Evaluation Rate analysis, demonstrating CorrSteer’s ability to minimize side effects while improving task performance.
Figure 5: SER comparison between different CorrSteer variants for Gemma 2 2B.
6 Discussion
This work establishes a viable and efficient approach for SAE-based steering, providing effective control across diverse applications. The interpretable feature combinations that yield performance improvements support the hypothesis that linear correlation serves as a meaningful unit for interpretable AI capabilities.
6.1 Feature Inspection
g A notable finding is that features selected by CorrSteer demonstrate task-relevant patterns that align with theoretical expectations Math-related features are consistently discovered across all tasks, proving beneficial even for bias mitigation and safety tasks. This universal correlation between mathematical features and accuracy aligns with DeepSeekMath (Shao et al., 2024)’s findings, where further pre-training on math-focused corpora yielded performance improvements across diverse tasks.
For BBQ features in LLaMA 3.1 8B, we observe:
While bias-related features were expected for the BBQ benchmark, these neutrality-focused features demonstrate high positive correlation. Conversely, explicit bias-related and choice-making features exhibit negative correlations:
These findings suggest that task-specific induced features contribute more to sample accuracy than meta-cognitive recognition features. Our ablation study further demonstrates that SAE-based sparse feature selection consistently outperforms raw activation steering across all evaluated tasks (Table 7).
6.2 Feature Set Transferability
The transferability of CorrSteer feature sets is evaluated across MMLU, MMLU-Pro, and BBQ benchmarks. Interestingly, our cross-task experiments reveal that MMLU features outperform task-specific features on BBQ Ambig and achieve comparable performance on MMLU-Pro, suggesting that some feature sets capture more generalizable reasoning patterns (Table 8). This transferability is attributed to their shared multiple-choice format, which requires similar structural feature patterns.
6.3 Task-Level Circuit
CorrSteer’s multi-layer approach relates to neural network circuit discovery research (Olah et al., 2020; Elhage et al., 2021). While emerging works focus on discovering task-specific circuits (Conmy et al., 2023; Marks et al., 2025; Ameisen et al., 2025; Lindsey et al., 2025; Sun, 2025), our steering vectors that work simultaneously across layers can be conceptualized as additive subgraphs of optimized task circuits, though they lack explicit interpretation of interactions and causality.
6.4 Side Effect Ratio
The primary challenge in AI steering concerns robustness for industrial applications, necessitating precise control mechanisms. Direct steering at each layer without updating original parameters based on token prediction distributions represents a key approach to minimize side effects. Theoretically, separating steering vectors across different activation spaces minimizes mutual interference in superpositioned states (Elhage et al., 2022).
6.5 Pooling Strategy Analysis
The results reveal interesting patterns across different pooling strategies (Table 5). Mean-pooling shows severe degradation on multi-token generation tasks (HarmBench: 0.00%, XSTest: 53.65%) where responses require multiple tokens. All-token pooling shows degraded performance on every task compared to max-pooling’s inference-time aggregation. This suggests that max-pooling better captures the critical activations needed for effective steering across all task types, while all-token pooling may introduce noise by including irrelevant token positions, and mean-pooling dilutes important signals by averaging across all tokens in longer sequences.
6.6 Computational Efficiency
Our streaming correlation implementation achieves O(1) memory complexity with respect to training set size, making the approach scalable to large datasets. The pipeline exhibits computational efficiency with minimal sample requirements (200-400) as demonstrated in Figure 4 and completes feature extraction within minutes. Static feature sets and coefficients at inference time eliminate SAE dependency during deployment.
6.7 Practical Applicability
CorrSteer operates as an auxiliary mechanism that captures correlations difficult to detect through supervised fine-tuning, applicable with minimal side effects even after fine-tuning. The automated pipeline enables rapid deployment across tasks and domains without hyperparameter tuning. The method shows effectiveness across two model families, as demonstrated on both Gemma-2 2B and LLaMA-3.1 8B models. Importantly, our approach has broader implications for AI safety, as demonstrated by its effectiveness in both bias mitigation and amplification (Table 9), highlighting the need for responsible deployment of such steering capabilities.
7 Conclusion
This work introduces CorrSteer, a fully automated pipeline for SAE-based language model steering that leverages correlation analysis. Our approach addresses key limitations of existing SAE steering methods by identifying task-relevant features without requiring manual feature exploration or contrastive datasets. Experimental validation across diverse benchmarks demonstrates CorrSteer’s effectiveness, consistently improving performance on question answering, bias mitigation, and safety evaluation tasks. Selected features for safety, mathematical, and refusal-related tasks reveal the underlying objectives and required capabilities that drive task performance.
Future Work
Several promising directions emerge from this work: Prompt Engineering Comparison: Future studies should compare CorrSteer with prompt engineering approaches, as prompt-based methods are expected to exhibit higher SER due to their less targeted intervention mechanisms. Dynamic Steering for Reasoning: The performance degradation observed in GSM8K reasoning tasks suggests the need for dynamic steering approaches that can adapt to the sequential nature of mathematical problem-solving, moving beyond static feature interventions. Orthogonal Feature Projection: To further minimize side effects, future work could explore feature filtering techniques that project out components already activated in baseline features before applying steering, potentially reducing interference with existing model capabilities.
Acknowledgments
The authors thank the teams behind Gemma Scope and LLaMA Scope for providing high-quality SAE releases that enabled this research. The authors also acknowledge Neuronpedia for providing automated feature descriptions that enhanced our interpretability analysis.
Limitations
The fundamental limitation of steering vectors is their static nature, which prevents adaptation to dynamic model behaviors. This constraint particularly affects reasoning tasks like GSM8K, where static steering cannot adequately handle the sequential nature of mathematical problem-solving. Our evaluation focuses primarily on discriminative and short-form generation tasks; long-form generation and creative tasks may require different approaches or modifications to our method.
Appendix A Appendix
A.1 Implementation Details
Feature Extraction:
Feature selection employs 4,000 samples across all datasets. For fair comparison, the same samples are used for training fine-tuning baselines. When datasets contain fewer than 4,000 samples, we use all available data. For datasets without predefined train/validation/test splits, we allocate 27% for training, 3% for validation, and 70% for testing. GSM8K uses 1,000 samples for feature selection with 50 samples reserved for validation.
Feature Steering:
Fine-tuning
Fine-tuning is performed using AdamW optimizer with learning rate 1e-5 (reduced to 5e-6 for small datasets <2000 samples), weight decay 0.01, and gradient clipping at norm 1.0. The training schedule includes 3% warmup steps followed by cosine annealing decay. Training proceeds for one epoch with 4,000 samples, using exact target supervision where prompt tokens are masked with -100 labels and only target spans contribute to the loss.
A.2 Generation Benchmark Results
A.3 Additional Results
Figure 6: Benchmark performance of CorrSteer variants compared with the default baseline on LLaMA 3.1 8B.
Task-Specific Analysis
MMLU: The global method selects features related to structured output formatting, addressing Gemma-2’s tendency to generate tokens outside the required A/B/C/D options. Post-steering, this hallucination issue is largely resolved.
MMLU-Pro: A similar issue occurs more severely due to the 10 options in MMLU-Pro. Constrained decoding, which samples tokens exclusively from available options, is applied to improve the model’s authentic capability, resulting in performance that remains higher than baseline, with CorrSteer-A achieving maximum performance.
Figure 7: SER comparison across datasets between different CorrSteer variants on LLaMA 3.1 8B.
@@ A.3 Additional Results
Figure 8: Frequency of activation samples across layers of Gemma 2 2B for SimpleQA (left) and HarmBench (right) tasks.

### v2

@@ 4 Experimental Setup
For safety benchmarks, both HarmBench (refusal) and XSTest (overrefusal) evaluate steering ability and contextual understanding.
Side Effect Evaluation. We measure Side Effect Ratio (SER) to quantify unintended performance degradations (Table 2). CorrSteer’s SER is compared against fine-tuning baselines across question-answering datasets. Additionally, we validate our positive-only feature selection by comparing performance when using negatively correlated features (Appendix 6). We also assess different pooling strategies to verify that inference-time token selection is optimal (Table 5).
Pooling Strategies for Feature Aggregation. To verify that our pooling design in Section 3.4 is robust, we conduct an ablation comparing three strategies for aggregating SAE activations across tokens: (i) mean-pooling, which averages activations across tokens; (ii) all-token pooling, which aggregates contributions from every position; and (iii) max-pooling, which selects the strongest activation. We evaluate these alternatives on GSM8K (reasoning), BBQ (bias), and HarmBench/XSTest (safety), covering both single-token and multi-token generation tasks. This setup isolates the effect of pooling and allows us to test whether CorrSteer’s empirically motivated default choices are consistently optimal across task types.
Feature Interpretability and Transferability Analysis. Performance-improving features are analyzed post-hoc using Neuronpedia descriptions to examine whether correlation-selected features exhibit semantic coherence (Appendix A.8.1). We analyze whether performance-improving features correspond to meaningful behaviors such as refusal, neutrality, or structured reasoning. Safe/unsafe tendency inspection and task-wise breakdowns test whether CorrSteer activates task-relevant semantics rather than spurious signals. Finally, we probe transferability by evaluating features selected on one benchmark (e.g., MMLU) on others (e.g., BBQ, MMLU-Pro) to test whether our method identifies generalizable circuits (Table 8).
5 Results and Discussion
Table 1 and Table 3 present comprehensive results across evaluation benchmarks. CorrSteer demonstrates improvements across question answering, bias mitigation, and safety benchmarks.
Table 1: Performance comparison across CorrSteer variants and other steering methods on Gemma-2 2B. Results are reported as mean ± standard deviation across 5 random seeds (3 for GSM8K). Within each method category, the best results are highlighted in bold, and the second-best results are highlighted in italics.
Method MMLU MMLU-Pro SimpleQA BBQ Ambig BBQ Disambig HarmBench XSTest GSM8K
CorrSteer Variants
Non-steered 52.21 ± 0.04 30.40 ± 0.21 3.78 ± 0.17 59.46 ± 0.21 75.38 ± 0.14 46.61 ± 2.78 86.35 ± 0.32 54.44 ± 0.35
CorrSteer-S 52.99 ± 0.47 30.38 ± 0.08 3.68 ± 0.07 62.39 ± 0.02 75.70 ± 0.01 46.61 ± 0.76 86.77 ± 0.48 53.63 ± 0.72
CorrSteer-P 54.70 ± 1.22 30.63 ± 0.13 3.80 ± 0.14 66.00 ± 2.15 76.48 ± 0.64 66.08 ± 20.20 86.46 ± 0.37 53.10 ± 0.74
CorrSteer-A 55.48 ± 0.59 30.93 ± 0.19 3.74 ± 0.07 62.06 ± 0.84 76.53 ± 0.23 73.75 ± 8.84 86.98 ± 1.45 40.34 ± 24.43
Other Methods
Fine-tuning 55.75 ± 0.09 35.32 ± 2.70 47.00 ± 0.33
SPARE (MI) 54.97 ± 0.87 30.84 ± 0.18 3.72 ± 0.04 64.81 ± 2.12 76.25 ± 0.59 65.43 ± 14.34 86.82 ± 0.76
DSG (Fisher) 52.81 ± 0.59 30.33 ± 0.16 3.66 ± 0.06 61.75 ± 1.39 75.61 ± 0.16 45.86 ± 1.76 86.35 ± 0.59
CAA 55.13 ± 1.00 28.01 ± 5.79 3.71 ± 0.07 62.40 ± 1.07 76.32 ± 0.40 43.14 ± 28.95 72.95 ± 17.50
5.1 Comparison with Baselines
Across benchmarks, CorrSteer-A and CorrSteer-P achieve the strongest results, with CorrSteer-P showing particular dominance in LLaMA-3.1 8B. This can be attributed to the less disentangled nature of LLaMA Scope features under superposition, which necessitates more aggressive pruning. Results on both Gemma-2 2B and LLaMA-3.1 8B confirm consistent improvement patterns.
The correlation-based approach consistently outperforms mutual information (MI) and Fisher information-based methods, supporting the faithfulness of SAE’s linear representation. This suggests that linear correlation-based feature extraction aligns with the linear latent space of SAEs, where features are designed to be linearly combined. Existing steering approaches rely on contrastive examples restricted to static contexts, while CorrSteer directly leveraging generation-time activations, extending SAE-based steering and achieving practical gains across QA, safety, and bias benchmarks.
Head-to-head comparison with CAA (Rimsky et al., 2024), DSG (Muhamed et al., 2025), or SPARE (Zhao et al., 2025) is not directly applicable since these methods require contrastive datasets rather than generation-time features. However, for comparison purposes, we apply our generation-time feature selection approach to these methods. For fair comparison, we applied the same test-time features and average positive coefficients across methods, with MI and Fisher information-based methods using substituted feature selection while CAA directly uses correct and incorrect answer activation differences. Furthermore, other methods also show improved performance when adapted to use generation-time features, demonstrating the effectiveness of our generation-time feature selection approach independent of the specific steering mechanism.
While fine-tuning achieves higher raw accuracy, CorrSteer offers advantages in side-effect reduction. On MMLU, CorrSteer-A achieves competitive accuracy (55.48% vs. 55.75%) while halving SER (0.20 vs. 0.41) (Table 1, Table 2). Although fine-tuning outperforms CorrSteer variants in raw accuracy on GSM8K and MMLU-Pro, CorrSteer maintains substantially lower SER across tasks. Moreover, CorrSteer can be layered on top of fine-tuned models as complementary enhancement.
Feature Collaboration and Circuit Effects. CorrSteer-A demonstrates superior performance in 5 out of 8 tasks, indicating that improvements often emerge from feature collaboration within circuits, even when individual feature steering yields limited benefit. Multi-layer approaches such as CorrSteer-A and CorrSteer-P consistently outperform the single-layer CorrSteer-S, aligning with prior findings on circuit-level interventions (Liu et al., 2024; Zhao et al., 2025).
Safety and Factuality. On HarmBench, selected features enhance refusal ability, achieving a 27.2% gain, though this primarily reflects increased refusal rather than fine-grained safety. In contrast, XSTest shows limited gains due to the benchmark’s over-refusal bias. This outcome is expected given the static nature of CorrSteer, which cannot easily separate benign from harmful requests. Similarly, on SimpleQA, CorrSteer yields only marginal improvement, confirming that the method enhances adherence to task requirements without introducing external factual knowledge. This is desirable, as it suggests CorrSteer modifies behavior rather than injecting content absent from the base model.
5.2 Efficiency and Scalability
CorrSteer serves as an auxiliary mechanism that identifies task-relevant features through generation-time correlations, complementing supervised fine-tuning and remaining effective when applied on top of fine-tuned models. The pipeline is fully automated, requires no hyperparameter tuning, and generalizes across tasks and domains with minimal adjustment. The streaming correlation algorithm operates with $\mathcal{O}(1)$ memory complexity relative to dataset size, ensuring scalability to large corpora. CorrSteer performs effectively with as few as 100 samples, though stable performance requires approximately 4,000 samples (Appendix A.1). Once steering vectors are extracted, inference requires no SAE dependency, since fixed feature sets and coefficients are sufficient.
Figure 3: Relation between sample counts and test performance, final matched count of selected features, and most correlated features from each Gemma-2 2B layer. Dotted lines show baseline default LLM performance and constrained decoding performance on MMLU answer options.
Training Sample Requirements:
As shown in Figure 3, CorrSteer performs effectively even with around 100 training samples, with no substantial improvements beyond 4,000 samples, making it practical for quick deployment. The high variance observed in CorrSteer-A for smaller datasets like GSM8K (1,000 samples) and HarmBench (108 samples) suggests that approximately 4,000 samples are recommended for stable performance.
5.3 Side Effect Trade-offs
To quantify trade-offs, we measure Side Effect Ratio (SER), capturing the proportion of harmful to beneficial side effects. Table 2 and Figure 4 show CorrSteer consistently lowers SER compared to fine-tuning while preserving accuracy and highlights CorrSteer-P as the most balanced strategy, while CorrSteer-S minimizes SER in safety-critical tasks such as HarmBench and BBQ.
Table 2: Side Effect Ratio (SER) results for CorrSteer variants on Gemma-2 2B across eight benchmarks. SER quantifies unintended side effects introduced by steering interventions, with values closer to 0 indicating stronger performance/safety preservation.
CorrSteer-S CorrSteer-P CorrSteer-A Fine-tuning
Task SER NEG POS SER NEG POS SER NEG POS SER NEG POS
MMLU 0.36 11 20 0.17 249 286 0.20 264 299 0.41 1108 1616
MMLU-Pro 0.42 8 11 0.42 30 41 0.42 39 54 0.46 357 418
GSM8K 0.56 20 16 0.67 31 15 0.52 63 59 0.65 213 116
BBQ Ambig 0.00 0 658 0.00 0 1532 0.08 53 649 - - -
BBQ Disambig 0.17 45 59 0.15 111 164 0.26 65 112 - - -
HarmBench 0.25 2 6 0.14 4 24 0.04 3 67 - - -
SimpleQA 0.17 1 5 0.19 3 13 0.35 6 11 - - -
XSTest 0.33 7 10 0.52 7 10 0.47 14 5 - - -
Figure 4: SER comparison between different CorrSteer variants for Gemma-2 2B.
5.4 Effect of Pooling Strategies on Steering
As discussed in Section 4, pooling determines how SAE activations are aggregated across tokens. To empirically validate these design choices, we conducted controlled experiments comparing mean-pooling, all-token pooling, and max-pooling across representative benchmarks covering reasoning (GSM8K), bias (BBQ), and safety (HarmBench, XSTest). The comparison is summarized in Table 5.
Our results reveal clear trends. On multi-token generation tasks, mean-pooling severely degrades performance (e.g., HarmBench: 0.00%, XSTest: 53.65%), confirming that averaging dilutes the sparse but informative signals needed for steering. All-token pooling similarly underperforms, suggesting that aggregating contributions from every token introduces substantial noise. By contrast, max-pooling consistently outperforms alternatives across tasks, capturing salient activations while filtering out irrelevant ones. These findings validate our choice of max-pooling as the default aggregation strategy for correlation-based feature selection and steering.
We note one exception: for long reasoning tasks such as GSM8K, mean-pooling provides more stable coefficient scaling, as max-pooling can produce excessively large coefficients when applied to every generated token. A more detailed ablation study, including per-task breakdowns and coefficient-scaling effects, is provided in Appendix A.4.
5.5 Feature Interpretability and Transferability
Selected features align with task requirements: structured output features dominate multiple-choice benchmarks (MMLU, BBQ), refusal-related features drive safety improvements (HarmBench), and domain-specific semantics contribute to specialized evaluations. Post-hoc analysis via Neuronpedia descriptions further supports their semantic relevance. Feature activation frequencies vary across tasks, with performance gains tracking activation dynamics (Appendix 7). Mathematical features also emerge across tasks, including bias and safety, consistent with findings that math-oriented pre-training improves broad accuracy (Shao et al., 2024).
For BBQ features in LLaMA-3.1 8B (full list in Appendix A.8.2), positively correlated features emphasize neutrality and balance:
These results suggest that task-specific semantic features contribute more to accuracy than general recognition features. Our ablation further confirms that SAE-based sparse feature selection outperforms raw activation steering (Table 7).
Feature Set Transferability. Cross-task experiments show that MMLU features transfer well, outperforming task-specific features on BBQ Ambig and performing comparably on MMLU-Pro (Table 8). This suggests that certain feature sets capture reasoning patterns shared by multiple-choice benchmarks.
Task-Level Circuit and Spurious Correlation. CorrSteer’s multi-layer steering relates to circuit discovery research (Olah et al., 2020; Elhage et al., 2021). While prior work isolates task-specific circuits (Conmy et al., 2023; Marks et al., 2025; Ameisen et al., 2025; Lindsey et al., 2025; Sun, 2025), our steering vectors act as additive subgraphs across layers. Restricting feature selection to activations during answer generation reduces spurious correlations, and interventions consistently improve performance (Table 1, Table 3), indicating the effectiveness of the selected feature sets.
Correlation for Selection, Intervention for Causality. CorrSteer employs correlation as a feature selection mechanism, then establishes causal relationships through direct steering interventions within the controlled LLM computational graph. Unlike spurious correlations with uncontrolled confounding variables, correlations within LLM circuits can be directly validated through residual stream intervention. The consistent performance improvements across tasks (Table 1, Table 3) demonstrate causal influence of selected features.
6 Conclusion and Limitations
This work introduces CorrSteer, a fully automated correlation-driven pipeline that enables generation-time discovery of steering-effective SAE features. By correlating task performance with specific activation patterns during inference, our method eliminates the dependency on contrastive datasets that has limited prior steering approaches. Across eight benchmarks, CorrSteer achieves substantial improvements in question answering, bias mitigation, and safety with minimal computational overhead and reduced side effects, while revealing semantically aligned steering circuits across multiple layers. By leveraging SAE’s inherently linear architecture where features are designed to be linearly combined, this design yields interpretable feature combinations without parameter modification, establishing linear correlation as a principled foundation for mechanistic interpretability.
Despite these advances, limitations remain. The fundamental constraint of steering vectors lies in their static nature, which prevents adaptation to dynamic model behaviors. This particularly affects tasks requiring contextual adaptation or multi-step reasoning, where static steering cannot adequately handle the conditional nature of problem-solving processes. Furthermore, our correlation-based approach exhibits increased performance variance with smaller sample sizes, and the task-optimized features show limited cross-task transferability beyond single-token generation scenarios.
Appendix A Appendix
A.1 Implementation Details
Feature Extraction:
Feature selection employs 4,000 samples across all datasets. For fair comparison, the same samples are used for training fine-tuning models. When datasets contain fewer than 4,000 samples, we use all available data. For datasets without predefined train/validation/test splits, we allocate 27% for training, 3% for validation, and 70% for testing. GSM8K uses 1,000 samples for feature selection with 50 samples reserved for validation.
Feature Steering:
@@ A.1 Implementation Details
Evaluation Metrics: For multiple-choice tasks (MMLU, MMLU-Pro, BBQ), exact match accuracy is used under zero-shot evaluation. All results are reported as mean ± standard deviation across multiple random seeds for statistical robustness: 5 seeds for most tasks, 3 seeds for GSM8K. For Gemma-2 2B, the non-steered MMLU performance (52.23%) is lower than the Gemma-2 2B-IT 5-shot result (56.1%) reported in the original Gemma paper due to the zero-shot setting and lack of in-context learning examples. For safety benchmarks, 1 - ASR (Attack Success Rate) is computed using a small refusal-detection language model. SimpleQA performance is measured using a small STS language model to match the expected answer, with more details in Appendix A.2.
@@
Fine-tuning
Fine-tuning hyperparameters are determined through empirical experimentation across tasks and dataset sizes. Fine-tuning is performed using AdamW optimizer with learning rate 1e-5 (reduced to 5e-6 for small datasets <2000 samples), weight decay 0.01, and gradient clipping at norm 1.0. The training schedule includes 3% warmup steps followed by cosine annealing decay. Training proceeds for one epoch with 4,000 samples, using exact target supervision where prompt tokens are masked with -100 labels and only target spans contribute to the loss.
A.2 Generation Benchmark Results
A.3 Additional Results
Figure 5: Benchmark performance of CorrSteer variants compared with the non-steered model on LLaMA-3.1 8B.
Table 3: Performance comparison between non-steered model and CorrSteer variants across BBQ, MMLU, MMLU-Pro, HarmBench, SimpleQA, and XSTest on LLaMA-3.1 8B. Results show accuracy (%) under zero-shot evaluation (single-shot for BBQ).
Task Non-steered Corrsteer-S CorrSteer-P CorrSteer-A
@@ A.3 Additional Results
Task-Specific Analysis
MMLU: The global method selects features related to structured output formatting, addressing Gemma-2 2B’s tendency to generate tokens outside the required A/B/C/D options. Post-steering, this hallucination issue is largely resolved.
MMLU-Pro: A similar issue occurs more severely due to the 10 options in MMLU-Pro. Constrained decoding, which samples tokens exclusively from available options, is applied to improve the model’s authentic capability, resulting in performance that remains higher than the non-steered model, with CorrSteer-A achieving maximum performance.
Figure 6: SER comparison across datasets between different CorrSteer variants on LLaMA-3.1 8B.
Table 4: Side Effect Ratio (SER) analysis for CorrSteer variants on LLaMA-3.1 8B across different benchmarks. SER values closer to 0 indicate better safety performance.
Corrsteer-S CorrSteer-P CorrSteer-A
@@
Figure 7: Frequency of activation samples across layers of Gemma-2 2B for SimpleQA (left) and HarmBench (right) tasks.